URide, a ride-hailing company, has operations worldwide and serves the US and EU markets through its URide West Unit and served APAC region through its URide East Unit. URide has no debt and the current market values of its two unit are $10.00bn and $2.00bn respectively. Gride, a ride-hailing company, is currently operating only in the APAC region, it has no debt and its current market value is $6.00bn. The equity ownership is provided in the table below. It is worth noticing that SBanc is a common shareholder in URide and GRide. Except for SBanc, none of the other URide shareholders has any stake in GRide and, again except for SBanc, none of the other GRide shareholders has any stake in URide.

                               Table 1. URide and GRide Ownership                                                                               

URide

GRide

SBanc

100.00 million shares

200.00 million shares

All but SBanc URide Shareholders

500.00 million shares

none

All but SBanc GRide Shareholders

none

600.00 million shares

Total Shares Outstanding

600.00 million shares

800.00 million shares

A market study indicates that the total market value in the APAC region is approximately $8.50bn. For now URide and GRide are the only two ride-hailing companies serving the APAC region. Seeing the additional 0.50bn synergies that could be created if one company only would serve the entire APAC market, URide and GRide decide to enter into an M&A transaction.

M&A - Partial Sale of a Subsidiary (continued) The terms of the transaction are as follow:

  • in light of the synergies of $0.50bn URide expects to receive in total (cash plus newly issued stock in GRide) a total of $2.00bn (as value of URide East Unit transferred to GRide) plus a premium calculated as a fraction of the synergies; the two parties agree that the synergies should be split according to the relative pre-merger values of URide East Unit and GRide (that is $2.00bn and $6.00bn);
  • GRide will issue new shares to be transferred to URide shareholders such that, post-merger, the stake of URide shareholders into GRide is 15.00%: if m is the number of new shares issued then the post-merger stake of URide shareholders into GRide is calculated as m / ( m + pre-merger GRide total number of shares );
  • GRide will pay cash C such that the sum of cash paid and the value of the m shares transferred to URide shareholders is equal to $2.00bn plus the premium required by URide;
  • URide shareholders do not surrender any of their URide shares, post-mergers they will still hold 100.00 million shares (SBanc) and 500 million shares (all but SBanc URide shareholders) and these shares indicate ownership of URide West Unit (which is technically the surviving URide company).

Question - Calculate the number of shares m that GRide will issue and the premium paid by GRide to URide.
Answer #1

m / (m + pre-merger GRide total number of shares ) = 15%

Hence, m / (m + 800) = 15% = 0.15

Hence, m = 0.15 x (m + 800) = 0.15m + 120

Hence, m = 120 / (1 - 0.15) =  141.18 million shares (Or 141,176,471 shares to be precise, if you have to answer in exact number of shares and not in million)

The two parties agree that the synergies should be split according to the relative pre-merger values of URide East Unit and GRide (that is $2.00bn and $6.00bn);

Hence, proportionate allocation of premium to URide East (as a fraction of the synergies) = 2 / ( 2 + 6) = 2 / 8 = 25%

Hence, Premium paid = Synergies valuation x 25% = 0.5 bn x 25% = $ 0.125 bn

(Please note, if premium is required to be entered into %age, please enter 25% as answer, if it's in $ bn, then enter $ 0.125 bn as your answer)